An obnoxious brat assigns his 0.0231-kilogram pet mouse Mickey the role of stunt mouse by stuffing him into a spring gun and shooting him up into the air at some angle. The gun\'s spring has a force constant of 84.7 N/m and is initially compressed 0.137 m. When Mickey reaches maximum height, he has a horizontal velocity with a speed of 2.05 m/s. How high above his initial location in the gun does Mickey soar? (In case you are worried, a miniature parachute opens at this point and Mickey lands safely. The brat is slightly less obnoxious than we thought.) Take g = 9.81 m/s2.
Answer #1

1) Energy from spring = 1/2 . k . x^2 = 0.5 . 84.7 . 0.137^2

(2) Energy that is still KE = 1/2 . m . v^2 = 0.5 . 0.0183 . 2.05^2

(3) The rest of the energy must be in GPE = m . g. h = 0.0231 . 9.81 . h

Take (2) from (1) and the remainder = (3)